# A hierarchical approach for multicontinuum homogenization in high contrast media

### v2

@@ Abstract
A recently developed upscaling technique, the multicontinuum homogenization method, has gained significant attention for its effectiveness in modeling complex multiscale systems. This method defines multiple continua based on distinct physical properties and solves a series of constrained cell problems to capture localized information for each continuum. However, solving all these cell problems on very fine grids at every macroscopic point is computationally expensive, which is a common limitation of most homogenization approaches for non-periodic problems. To address this challenge, we propose a hierarchical multicontinuum homogenization framework. The core idea is to define hierarchical macroscopic points and solve the constrained problems on grids of varying resolutions. We assume that the local solutions can be represented as a combination of a linear interpolation of local solutions from preceding levels and an additional correction term. This combination is substituted into the original constrained problems, and the correction term is resolved using finite element (FE) grids of varying sizes, depending on the level of the macropoint. By normalizing the computational cost of fully resolving the local problem to $\mathcal{O}(1)$, we establish that our approach incurs a cost of $\mathcal{O}(L\eta^{(1-L)d})$, highlighting substantial computational savings across hierarchical layers $L$, coarsening factor $\eta$, and spatial dimension $d$. Numerical experiments validate the effectiveness of the proposed method in media with slowly varying properties, underscoring its potential for efficient multiscale modeling.
@@ 1 Introduction
Among the various approaches developed for multiscale problems, an important class focuses on constructing local basis functions on fine grids and solving the resulting algebraic systems on coarse grids. This approach balances computational efficiency with the preservation of fine-scale details. Key methods in this category include the localized orthogonal decomposition method [23, 17], which generates localized multiscale basis functions through systematic decomposition of solution spaces, effectively capturing fine-scale information while reducing computational cost. The multiscale finite element method (MsFEM) [21, 20, 30, 22] employs precomputed basis functions to account for fine-scale heterogeneities, enabling accurate coarse-scale simulations. The generalized multiscale finite element method (GMsFEM) [13, 6, 2, 27, 11, 31, 16] extends the multiscale finite element method (MsFEM) by introducing symmetric and adaptive basis function construction. The constraint energy minimizing generalized multiscale finite element method (CEM-GMsFEM) [8, 29, 25, 28] further constructs local basis functions by solving energy minimization problems within oversampled domains.
@@
In contrast, some problems necessitate a focus solely on macroscopic variables, circumventing the need to resolve fine-scale details explicitly. Upscaling techniques establish a connection between microscopic and macroscopic scales by assuming that microscale solutions can be represented using different ansatzes of the macroscale solutions. Examples include homogenization [19, 12] and the heterogeneous multiscale method [1, 18]. In this paper, we focus on the multicontinuum homogenization method [7, 14, 26, 15, 3]. This method defines multiple continua to represent distinct physical regions. The macroscopic coefficient is obtained by solving a series of constraint problems in an oversampled representative volume element (or target coarse block) region. The constraints in the cell problems are used to represent different behaviors for different continua, which has been widely investigated in nonlocal multicontinuum upscaling (NLMC) methods [10, 9], as well as in the CEM-GMsFEM [8]. In fact, the accuracy of the upscaling coefficient depends heavily on the finite element (FE) grid used to solve the cell problems. In practice, since local geometries or material properties are typically non-periodic, resolving all cell problems on very fine grids is computationally expensive. For this reason, we introduce a general framework to hierarchically compute the upscaling coefficient.
@@
In this work, we propose a general framework for hierarchical multicontinuum homogenization. Our approach mainly consists of three steps: First, we divide the macropoints (the centers of the coarse blocks) into a hierarchical grid, similar to previous works [24, 4, 5]. Then, we define the FE spaces for different levels. Finally, we construct the local solutions inductively for different levels. We assume that the local solutions can be expressed as a combination of the linear interpolation of local solutions from other macropoints at preceding levels and an additional correction term. After substituting this combination into the original constraint problem, we obtain a constraint problem for the correction term. In this way, the computational cost mainly stems from numerically solving the correction term. By using the nested FE spaces we define, the computational cost can be reduced compared to using the same fine FE grid for each macropoint. We also discuss the detailed computational cost of the original multicontinuum homogenization method and the hierarchical multicontinuum homogenization method. In addition, we design four numerical examples to test the performance of our method. We compute three types of errors: (1) the error from the original multicontinuum homogenization method, (2) the error of the hierarchical multicontinuum homogenization method, and (3) the error of the hierarchical part. Our numerical results show that if the physical properties do not vary drastically, our approach achieves high accuracy. Even when the physical regions have dramatic variations, it can still provide acceptable accuracy.
@@ 2 Preliminaries
In this section, we present the model problem and review the multicontinuum homogenization framework introduced in [14, 7]. Let $\Omega\subset\mathbb{R}^{d}$ ($d=2,3$) be a convex region. We consider the following equation:
@@
where $V$ is $H_{0}^{1}(\Omega)$. In this paper, we use $h$ to represent the mesh size that can resolve the heterogeneities caused by the coefficients or geometries and achieve high accuracy. To avoid the unaffordable algebra system from that mesh size, we use the multicontinuum homogenization method to build an upscaling model and solve it at the macroscopic scale.
@@
We present the computational grid for multicontinuum homogenization. In Figure 1, the computation domain $\Omega$ is divided into a series of coarse blocks. We select a coarse block $K_{p}$, whose center is $x_{p}^{*}$. In the rest of this paper, the superscript * is used to denote the macropoint. The representative volume element (RVE) $R_{p}$ is used to capture the local heterogeneity information at the macropoint $x_{p}^{*}$. To reduce the influence of artificial boundary conditions, we apply oversampling to $R_{p}$ to get $R_{p}^{+}$.
The continuum is defined through different characteristic functions $\psi_{i}^{p}$ in the corresponding representative volume element $R_{p}$. Specifically, $\psi_{i}^{p}=1$ in continuum $i$ and 00 elsewhere. We use two continua for illustration, but this approach can be extended to multiple continua. We then assume that the unknown variable $u$ can be expanded as a series of products of macroscopic variables $U_{i}$ and local microscopic functions $\phi$’s:
@@
In general, we only consider the constant and linear behavior of macroscopic functions. That is, we assume $u\approx\phi_{i}U_{i}+\phi_{i}^{m}\partial_{m}U_{i}$. The local solutions $\phi$ are constructed by numerically solving a constrained problem in an oversampled domain $R_{p}^{+}$. The first-type saddle point problem is formulated by imposing constraints designed to capture the macroscopic average behavior of multiple continua.
$\int_{R_{p}^{+}}\kappa\nabla\phi_{i}\cdot v-\sum_{j,q}\frac{\beta_{ij}^{q}}{\int_{R_{q}}\psi_{j}}\int_{R_{q}}\psi_{j}v=0,$ (3)
@@
$\int_{R_{p}^{+}}\kappa\nabla\phi_{i}\cdot v-\sum_{j,q}\frac{\beta_{ij}^{q}}{\int_{R_{q}}\psi_{j}}\int_{R_{q}}\psi_{j}v=0,$ (4)
@@
$\int_{R_{p}}\kappa\nabla(\phi_{i}U_{i})\nabla v\approx U_{i}(x_{p}^{*})\int_{R_{p}}\kappa\nabla\phi_{i}\nabla v,\quad\int_{R_{p}}\kappa\nabla(\phi_{i}^{m}\nabla_{m}U_{i})\nabla v\approx\nabla_{m}U_{i}(x_{p}^{*})\int_{R_{p}}\kappa\nabla\phi_{i}^{m}\nabla v.$ (5)
@@
$U_{i}V_{j}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\kappa\nabla\phi_{i}\nabla\phi_{j}+V_{j}\nabla_{m}U_{i}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\kappa\nabla\phi_{i}^{m}\nabla\phi_{j}+U_{i}\nabla_{n}V_{j}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\nabla\phi_{i}\nabla\phi_{j}^{n}$
$+\nabla_{m}U_{i}\nabla_{n}V_{j}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\kappa\nabla\phi_{i}^{m}\nabla\phi_{j}^{n}=\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}f\phi_{j},$
@@ 3 Hierarchical multicontinuum homogenization
The main goal of hierarchical multicontinuum homogenization is to use a different FE grid size to solve the cell problems, while ensuring that all of them achieve the finest grid accuracy. All of the local solutions can be summarized as the following abstract problem:
$\Phi_{p}=\min_{v\in V(R_{p})}\mathcal{A}^{\epsilon}(v)\quad\textnormal{subjectto some constraints.}$. (8)
The constraints are detailed in Eqs. (3) and (4). Hereafter, $\Phi_{p}$ is used to represent the local solutions at macropoint $x_{p}^{*}$, regardless of the specific type of local solutions. Since solving the cell problems for every macroscopic point using a very fine mesh remains a huge challenge, we design a hierarchical multicontinuum homogenization method to overcome this difficulty. For a given macroscopic point $x_{p}^{*}$, we assume that the local solutions $\Phi_{p}$ can be expressed as
$\Phi_{p}=\Xi_{p}+\mathcal{I}_{p}(\Phi).$ (9)
Here, $\mathcal{I}_{p}$ is a linear interpolation operator at $x_{p}^{*}$ using the local solutions that have already been constructed, and $\Xi_{p}$ is a correction term solved using a different size of the fine grid mesh. The computational process of our algorithm involves three steps: First, we build a dense hierarchical macrogrid $\{\mathcal{S}_{n}\}_{n=1}^{L}$ for the macropoint set $\mathcal{T}$. Second, we define a nested FE space $\{V_{n}\}_{n=1}^{L}$ to matching the hierarchical macrogrid $\{\mathcal{S}_{n}\}_{n=1}^{L}$. Finally, we substitute Eq. (9) into the constraint problem (8) and solve for the correction term $\Xi_{p}$ within the corresponding FE space. The framework of our method is listed in Algorithm 1.
@@
In the following, we focus on the details of hierarchically computing the local solutions.
@@ 3.1 Build hierarchy of macrogrids
A hierarchical macrogrid can help us judiciously decide the FE mesh size for cell problems at different macrogrid points. We first build a nested macroscopic grid (cf. [4, 24, 5]) for the macropoint set $\mathcal{T}$, denoted as
@@ 3.3 Calculating the correction term
where $\phi_{i,t}$ represents the average behavior of the $i$-th continuum at the neighboring macropoint $x_{t}^{*}$, and $\phi_{i,t}^{m}$ represents the linear behavior of the $i$-th continuum along the $m$-th axis at the neighboring macropoint $x_{t}^{*}$. For $x_{p}^{\in}\mathcal{S}_{1}$, as no pre-constructed level exists, we define $\mathcal{I}_{p}=0$. Note that the linear interpolation coefficients satisfy $\sum_{t=1}^{r}c_{t}=1$. For the macropoint $x_{p}^{*}\in\mathcal{S}_{n}$, we substitute the local solutions $\phi_{i,p}$ from (11) into (3), and we obtain:
@@
In Figure 4, we color the coarse blocks to illustrate the 3-level hierarchical macrogrids and show the fine grid structure for the cell problems. We use blue to represent the first level $\mathcal{S}_{1}$, brown for the second level $\mathcal{S}_{2}$, and light gray for $\mathcal{S}_{3}$. In our numerical experiments, we only consider the first level for the interpolation operator, i.e., $\mathcal{I}_{p}=\Phi_{t}$ with $x_{t}^{*}\in\mathcal{S}_{1}$. More sophisticated combinations can indeed be used. From Figure 4, we can clearly observe that the number of first-level macropoints is the smallest, while the third level has the most. This highlights that we perform the finest computations on a few macropoints, but coarser computations on the majority of macropoints.
[…]

### v6

@@ Abstract
A recently developed upscaling technique, the multicontinuum homogenization method, has gained significant attention for its effectiveness in modeling complex multiscale systems. This method defines multiple continua based on distinct physical properties and solves a series of constrained cell problems to capture localized information for each continuum. However, solving all these cell problems on very fine grids at every macroscopic point is computationally expensive, which is a common limitation of most homogenization approaches for non-periodic problems. To address this challenge, we propose a hierarchical multicontinuum homogenization framework. The core idea is to define hierarchical macroscopic points and solve the constrained problems on grids of varying resolutions. The local solutions are decomposed into the linear interpolation of contributions inherited from preceding levels and an additional correction term. This combination is substituted into the original constrained problems, and the correction term is resolved using finite element (FE) grids of varying sizes, depending on the level of the macropoint. By normalizing the computational cost of fully resolving the local problem to $\mathcal{O}(1)$, we establish that our approach incurs a cost of $\mathcal{O}(L\eta^{(1-L)d})$, highlighting substantial computational savings across hierarchical layers $L$, coarsening factor $\eta$, and spatial dimension $d$. Numerical experiments validate the effectiveness of the proposed method in media with slowly varying properties, underscoring its potential for efficient multiscale modeling.
@@ 1 Introduction
Among the various approaches developed for multiscale problems, an important class focuses on constructing local basis functions on fine grids and solving the resulting algebraic systems on coarse grids. This approach balances computational efficiency with the preservation of fine-scale details. Key methods in this category include the localized orthogonal decomposition method [maalqvist2014localization, henning2014localized], which generates localized multiscale basis functions through systematic decomposition of solution spaces, effectively capturing fine-scale information while reducing computational cost. The multiscale finite element method (MsFEM) [hou1997multiscale, hou1999convergence, ye2020convergence, jiang2017reduced, huang2001partition] employs precomputed basis functions to account for fine-scale heterogeneities, enabling accurate coarse-scale simulations. The generalized multiscale finite element method (GMsFEM) [efendiev2013generalized, calo2016randomized, alikhanov2025multiscale, xie2025time, xie2025multiscale, chung2014adaptive, ye2024robust, fu2024efficient] extends the multiscale finite element method (MsFEM) by introducing symmetric and adaptive basis function construction. The constraint energy minimizing generalized multiscale finite element method (CEM-GMsFEM) [chung2018constraint, ye2023constraint, wang2024multiscale, xie2024cem] further constructs local basis functions by solving energy minimization problems within oversampled domains, and exhibits convergence with respect to the coarse mesh size.
@@
In contrast, some problems necessitate a focus solely on macroscopic variables, circumventing the need to resolve fine-scale details explicitly. Upscaling techniques establish a connection between microscopic and macroscopic scales by assuming that microscale solutions can be represented using different ansatzes of the macroscale solutions. Examples include homogenization [hornung2012homogenization, cioranescu1999introduction] and the heterogeneous multiscale method [abdulle2012heterogeneous, henning2009heterogeneous]. In this paper, we focus on the multicontinuum homogenization method [chung2024multicontinuum, efendiev2023multicontinuum, xie2025multicontinuum, efendiev2025multicontinuum, bai2026multicontinuum]. This method defines multiple continua to represent distinct physical regions. The macroscopic coefficient is obtained by solving a series of constraint problems in an oversampled representative volume element (or target coarse block) region. The constraints in the cell problems are used to represent different behaviors for different continua, which has been widely investigated in nonlocal multicontinuum upscaling (NLMC) methods [chung2018non, chung2021nonlocal], as well as in the CEM-GMsFEM [chung2018constraint]. In fact, the accuracy of the upscaling coefficient depends heavily on the finite element (FE) grid used to solve the cell problems. In practice, since local geometries or material properties are typically non-periodic, resolving all cell problems on very fine grids is computationally expensive. This motivates the development of a hierarchical strategy for computing the upscaling coefficient.
@@
In this work, we propose a hierarchical multicontinuum homogenization framework. The method consists of three main components. First, the macropoints (centers of coarse blocks) are organized into a hierarchy of grids. Next, distinct finite element spaces are assigned to different levels. Finally, the local solutions are constructed inductively, with each solution expressed as a combination of a linear interpolation from macropoints at preceding levels and a correction term. Substituting this representation into the original constrained problem yields a reduced problem for the correction term, which becomes the main computational task. Since the correction problems are solved in nested FE spaces, the overall cost is significantly lower than solving all problems on the same fine grid. We also provide a detailed cost comparison between the original and hierarchical formulations. Numerical experiments in two and three dimensions confirm the efficiency and accuracy of the proposed framework: when physical properties vary moderately, the hierarchical approach achieves high accuracy, even in cases with dramatic variations, the results remain acceptable.
Over the past decades, hierarchical ideas [park2020hierarchical, brown2013efficient, brown2017hierarchical] and multilevel methods [briggs2000multigrid, xu2017algebraic, cheng2022scalable, efendiev2015multilevelcicp, li2018multilevel, ye2025highly, efendiev2015multilevelmms] have been extensively developed and widely applied in large-scale scientific computing, including reservoir simulation and porous media flow. Most of these approaches, however, either explicitly resolve fine-grid solutions or rely on scale-separation assumptions to construct efficient macroscopic models or to establish convergence analyses. In contrast, the main contribution of this work is an efficient upscaling technique that bypasses direct fine-grid computation while producing a computable and accurate macroscopic model, without assuming scale separation or periodicity of the permeability field. Our approach involves only two levels. The first level is used to construct the computable macroscopic model, while the second level is employed to compute an efficient macroscopic coefficient. To reduce the computational cost at the second level, a hierarchical algorithm is applied. To the best of our knowledge, this is also the first work to present three-dimensional numerical examples for the multicontinuum homogenization method.
@@ 2 Preliminaries
In this section, we present the model problem and review the multicontinuum homogenization framework introduced in [efendiev2023multicontinuum, chung2024multicontinuum]. Let $\Omega\subset\mathbb{R}^{d}$ ($d=2,3$) be a convex region. We consider the following equation:
@@
where $V$ is $H_{0}^{1}(\Omega)$. We use $h$ to denote the mesh size required to resolve the heterogeneities induced by the coefficients or geometries and to achieve high accuracy. The fine-scale finite element partition is assumed to be general and may consist of either structured or unstructured meshes, provided it forms a shape-regular family. For clarity of presentation, we illustrate the method using rectangular elements. To avoid the unaffordable algebra system from that mesh size, we use the multicontinuum homogenization method to build an upscaling model and solve it at the macroscopic scale.
@@
We present the computational grid used for multicontinuum homogenization. The computational domain $\Omega$ is partitioned into a series of coarse blocks, and we denote the center of a coarse block $K_{p}$ by $x_{p}^{*}$. The coarse partition is in principle arbitrary and may be adapted to the heterogeneities of the operator $\mathcal{A}^{\epsilon}$. In the examples below, we employ rectangular coarse blocks solely for ease of exposition. In the following, the superscript * is used to indicate a macropoint. To capture the local heterogeneity at each macropoint $x_{p}^{*}$, a representative volume element (RVE) $R_{p}$ is defined. To mitigate the influence of artificial boundary conditions, the RVE is extended via oversampling to $R_{p}^{+}$. This procedure is particularly necessary in high-contrast cases, and the number of oversampling layers depends on the coarse mesh size and the contrast of the medium, as analyzed in [maalqvist2014localization, chung2018constraint].
The continuum is defined through different characteristic functions $\psi_{i}^{p}$ in the corresponding representative volume element $R_{p}$. Specifically, $\psi_{i}^{p}=1$ in continuum $i$ and 0 elsewhere. We use two continua for illustration, but this approach can be extended to multiple continua. We then assume that the unknown variable $u$ can be expanded as a series of products of macroscopic variables $U_{i}$ and local microscopic functions $\phi$’s:
@@
The formulation of the local solutions $\phi$ is inspired by [chung2018constraint, chung2021nonlocal], where they are cast as constrained cell problems. Specifically, the local solutions are obtained by minimizing the energy associated with the PDE operator $\mathcal{A}^{\epsilon}$ in a local domain $R_{p}^{+}$, subject to constraints that preserve certain integral properties of the macroscopic functions. In this work, we restrict our attention to local problems corresponding to constant and linear macroscopic behaviors. In particular, we approximate $u$ as $u\approx\phi_{i}U_{i}+\phi_{i}^{m}\partial_{m}U_{i}$.
The first-type saddle point problem is formulated by imposing constraints designed to capture the macroscopic average behavior of multiple continua.
$\int_{R_{p}^{+}}\kappa\nabla\phi_{i}\cdot\nabla v-\sum_{j,q}\frac{\beta_{ij}^{q}}{\int_{R_{q}}\psi_{j}}\int_{R_{q}}\psi_{j}v=0,$ (3)
@@
$\int_{R_{p}^{+}}\kappa\nabla\phi_{i}\cdot\nabla v-\sum_{j,q}\frac{\beta_{ij}^{q}}{\int_{R_{q}}\psi_{j}}\int_{R_{q}}\psi_{j}v=0,$ (4)
@@
$\int_{R_{p}}\kappa\nabla(\phi_{i}U_{i})\nabla v\approx U_{i}(x_{p}^{*})\int_{R_{p}}\kappa\nabla\phi_{i}\nabla v,$ (5)
$\int_{R_{p}}\kappa\nabla(\phi_{i}^{m}\nabla_{m}U_{i})\nabla v\approx\nabla_{m}U_{i}(x_{p}^{*})\int_{R_{p}}\kappa\nabla\phi_{i}^{m}\nabla v.$
@@
$\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}f\phi_{j}=U_{i}V_{j}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\kappa\nabla\phi_{i}\nabla\phi_{j}+V_{j}\nabla_{m}U_{i}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\kappa\nabla\phi_{i}^{m}\nabla\phi_{j}$
$+U_{i}\nabla_{n}V_{j}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\nabla\phi_{i}\nabla\phi_{j}^{n}+\nabla_{m}U_{i}\nabla_{n}V_{j}\sum_{p}\frac{|K_{p}|}{|R_{p}|}\int_{R_{p}}\kappa\nabla\phi_{i}^{m}\nabla\phi_{j}^{n},$
@@ 3 Hierarchical multicontinuum homogenization
The main goal of hierarchical multicontinuum homogenization is to use a different FE grid size to solve the cell problems, while ensuring that all of them achieve the finest grid accuracy. The local solutions (3) and (4) can be summarized as follows:
$\Phi_{p}=\arg\min\limits_{v\in V(R_{p}^{+})}\mathcal{J}^{\epsilon}(v),\quad\textnormal{s.t. constraints,}$ (8)
where $\mathcal{J}^{\epsilon}(v)=\tfrac{1}{2}\int_{R_{p}^{+}}\kappa|\nabla v|^{2}$. The constraints are detailed in Eqs. (3) and (4). Hereafter, $\Phi_{p}$ is used to represent the local solutions at macropoint $x_{p}^{*}$, regardless of the specific type of local solutions. Since solving the cell problems for every macroscopic point using a very fine mesh remains a huge challenge, we design a hierarchical multicontinuum homogenization method to overcome this difficulty.
For each macropoint $x_{p}^{*}$, the local solution $\Phi_{p}$ is represented as a decomposition into two components: (i) an interpolation term $\mathcal{I}_{p}(\Phi)$, which resolves the cell problems on a very fine mesh inherited from preceding levels, and (ii) a correction term $\Xi_{p}$, computed on a coarser mesh to account for local variations. This decomposition reads
$\Phi_{p}=\mathcal{I}_{p}(\Phi)+\Xi_{p}.$ (9)
The computational process of our algorithm involves three steps: First, we build a dense hierarchical macrogrid $\{\mathcal{S}_{n}\}_{n=1}^{L}$ for the macropoint set $\mathcal{T}$. Second, we define a nested FE space $\{V_{n}\}_{n=1}^{L}$ to match the hierarchical macrogrid $\{\mathcal{S}_{n}\}_{n=1}^{L}$. Finally, we substitute Eq. (9) into the constraint problem (8) and solve for the correction term $\Xi_{p}$ within the corresponding FE space. The framework of our method is summarized in Algorithm 1.
In the following, we present the details of the hierarchical computation of the local solution.
@@ 3.1 Build hierarchy of macrogrids
A hierarchical macrogrid can help us judiciously decide the FE mesh size for cell problems at different macrogrid points. We first build a nested macroscopic grid (cf. [brown2013efficient, park2020hierarchical, brown2017hierarchical]) for the macropoint set $\mathcal{T}$, denoted as
@@ 3.3 Calculating the correction term
where $\phi_{i,t}$ represents the average behavior of the $i$-th continuum at the neighboring macropoint $x_{t}^{*}$, and $\phi_{i,t}^{m}$ represents the linear behavior of the $i$-th continuum along the $m$-th axis at the neighboring macropoint $x_{t}^{*}$. For $x_{p}^{*}\in\mathcal{S}_{1}$, as no pre-constructed level exists, we define $\mathcal{I}_{p}=0$. Note that the linear interpolation coefficients satisfy $\sum_{t=1}^{r}c_{t}=1$. For the macropoint $x_{p}^{*}\in\mathcal{S}_{n}$, we substitute the local solutions $\phi_{i,p}$ from (11) into (3), and we obtain:
@@
- Solving by $h$- Solving by $2h$- Solving by $4h$
@@
In Figure 4, the coarse blocks are color-coded to illustrate the three-level hierarchical macrogrids. The color scheme is as follows: blue for the first level $\mathcal{S}_{1}$, brown for the second level $\mathcal{S}_{2}$, and light gray for the third level $\mathcal{S}_{3}$. The figure also shows the FE grids used to resolve the cell problems for constructing the correction terms (12) and (13). As observed in Figure 4, the number of macropoints is smallest at the first level and largest at the third, reflecting the strategy of performing the most expensive fine-scale computations on a limited number of macropoints while relying on coarser approximations for the majority.
In the numerical experiments, only macropoints at the first level are employed in the interpolation operator, i.e., $\mathcal{I}_{p}=\Phi_{t}$ with $x_{t}^{*}\in\mathcal{S}_{1}$, although more sophisticated combinations can be considered. Figure 4 highlights a first-level macropoint along with neighboring points that inherit its highly resolved local solutions. We observe that only one macropoint requires solving the cell problem on a grid of size $\mathcal{O}(h^{-2})$, three macropoints on $\mathcal{O}(4h^{-2})$, and twelve on $\mathcal{O}(16h^{-2})$. Compared with fully resolving all sixteen macropoints at the finest level $\mathcal{O}(h^{-2})$, this hierarchical strategy achieves significant computational savings. More general computational costs are discussed in the next subsection, and the performance and cost reductions associated with different macrogrid levels are reported in the numerical examples section.